OpenStudy (anonymous):

Determine algebraically where the cubic polynomial that has zeros at 2, 3, and -5 and passes through the point (4,36) has a value of 120?

OpenStudy (anonymous):

y=a(x-2)(x-3)(x+5) 36=a(2)(1)(9) 36=18a a =2. So y=2(x-2)(x-3)(x+5)

OpenStudy (anonymous):

What about the value of 120?